OpenStudy (maeganr):

Solve the following triangles for all sides and angles: a=5 b=8 C=40 degrees, what is A, B, and c

OpenStudy (wolf1728):

Well it would help to see a graphic:

OpenStudy (wolf1728):

When you know two sides and the included angle, then you use the Law of Cosines to solve the other side and other 2 angles.

OpenStudy (maeganr):

I know all of that I just can't figure out how to make it work to get A B and c

OpenStudy (wolf1728):

Here's how to make it work: Let's use the 3rd formula in the graphic: c^2 = a^2 +b^2 - 2ab * cosine C c^2 = 5^2 + 8^2 -2*5*8 * cosine (40) Can you solve for side c^2?

OpenStudy (wolf1728):

okay a = 5 What is a^2?

OpenStudy (maeganr):

27.7164

OpenStudy (maeganr):

I got 27.7164 but the answer key says it's 5.3 ?

OpenStudy (wolf1728):

You have correctly calculated c^2 Now what does c equal?

OpenStudy (maeganr):

13.8582

OpenStudy (wolf1728):

If c^2 = 27.7164 then what is c?

OpenStudy (maeganr):

13.8582

OpenStudy (wolf1728):

c is the square root of 27.7164 which is?

OpenStudy (maeganr):

5.2646

OpenStudy (wolf1728):

Correct!

OpenStudy (maeganr):

what's A and B then

OpenStudy (wolf1728):

If you know all 3 sides (which we now do) we can use the Law of Cosines again.

OpenStudy (wolf1728):

cos A = (b^2 + c^2 - a^2) / 2*b*c

OpenStudy (wolf1728):

cos A = (8^2 + 5.2646^2 -5^2) / (2 * 8 * 5.2646) Can you calculate angle A?

OpenStudy (maeganr):

I got 1404.9324

OpenStudy (wolf1728):

cos A = (64 + 17.716 -25) / 84.2336 cos A = 56.716 / 84.2336 cos A = 0.67332 Do you know how to find the arc cosine of A?

OpenStudy (maeganr):

wheres the 17.716 from

OpenStudy (wolf1728):

My mistake that should be (8^2 + 5.2646^2 -5^2) / (2 * 8 * 5.2646) = (64 + 27.716 - 25) / 84.2336 = 66.716 / 84.2336 =0.7920354823 can you find the arc cosine of 0.7920354823?

OpenStudy (wolf1728):

Here's a good online trig calculator http://www.1728.org/trigcalc.htm Now can you calculate the arc cosine of .7920354823?

OpenStudy (wolf1728):

Did you go to the online trig calculator?

OpenStudy (maeganr):

i got 37.6238 idk

OpenStudy (wolf1728):

That is correct! Now (and finally) the LAST thing to do is calculate the third angle which is quite easy

OpenStudy (maeganr):

102.3762

OpenStudy (wolf1728):

Precisely!! Now you see how it is done!

OpenStudy (maeganr):

thank you

OpenStudy (wolf1728):

u r welcome :-)